OpenStudy (anonymous):

Does anyone know how to integrate (1-(sinx)^3)/(sinx)^2? I already tried a few things, but everytime i get stuck...

OpenStudy (helder_edwin):

i think u can work on this \[\large \int\frac{1-\sin^3x}{\sin^2x}\,dx= \int\left(\frac{1}{\sin^2x}-\frac{\sin^3x}{\sin^2x}\right)\,dx \]

OpenStudy (helder_edwin):

\[\large =\int\left(\frac{1}{\sin^2x}-\sin x\right)\,dx \]

OpenStudy (helder_edwin):

u can use tables for this